OpenStudy (australopithecus):

find a value of n such that sn is within 0.0004 \[\sum_{n=1}^{\infty} \frac{2}{n^{5}}\]

OpenStudy (amistre64):

this reminds me of a remainder function thingy

OpenStudy (amistre64):

\[S=S_n+R_n\]\[R_n < .0004\]if memory serves

OpenStudy (amistre64):

\[R_n: 2\int_n^{\infty}x^{-5}dx<.0004\] \[\int_n^{\infty}x^{-5}dx<.0002\] \[-\frac{1}{4}\infty ^{-4}+\frac{1}{4}n^{-4}<.0002\] \[\frac{1}{4}n^{-4}<.0002\] \[n^{-4}<.0008\] \[n^{4}>\frac{1}{.0008}\] \[n>\sqrt[4]{\frac{1}{.0008}}\] maybe

OpenStudy (amistre64):

n > 5.95 so id chk at n=5 and n=6 to be safe

OpenStudy (amistre64):

to inf = x.0738555... to 5 = x.0733235... to 6 = x.0735807... .0738555 .0735807 ---------- .0002748 .0738555... .0733235... ----------- .0005320 id say n=6 wins

OpenStudy (australopithecus):

how did you get the values: to inf = x.0738555... to 5 = x.0733235... to 6 = x.0735807...

OpenStudy (amistre64):

i plugged the summation into wolframalpha.com to make the dbl chk easier for me

OpenStudy (australopithecus):

oh I see thanks I must have been typing in my problem incorrectly

OpenStudy (amistre64):

sum (from a to b) f(n) i think is what i used

OpenStudy (australopithecus):

so you are looking for the value of n that is closest to 0.0004?

OpenStudy (amistre64):

you are looking for the value of n that insures that you are no more than .0004 in error

OpenStudy (amistre64):

the smallest n that is

OpenStudy (amistre64):

(sum at infinity) - (sum at n) is less than .0004

OpenStudy (amistre64):

since: \[S_{\infty}=S_n+R_n \]then\[S_{\infty}-S_n=R_n \] but we want it such that \[S_{\infty}-S_n<.0004\]which means that we need to find a suitable \[R_n < .0004\]

OpenStudy (amistre64):

the remainder part is going to fall someplace between \(\int R_n\) and \(\int R_{n+1}\)

OpenStudy (amistre64):

\[\int_{n}^{\infty}f(x)dx\ \le \sum_{n}^{\infty}f(n)\le\ \int_{n+1}^{\infty}f(x)dx\] or something like that

OpenStudy (amistre64):

i prolly got my n and n+1 mixed around tho

OpenStudy (australopithecus):

I see thanks